The number of accidents per week at a hazardous intersection varies with mean 2.2 and standard deviation 1.4 accidents. This variable is discrete. It is certainly not normaly distributed. Let x-bar be the mean number of accidents per week at the intersection during a year.(52 weeks).
Question: What is the approximate chance that there are fewer than 130 accidents at the intersection a year? (Restate in terms of x-bar)

Answer by stanbon(75887)

The number of accidents per week at a hazardous intersection varies with mean 2.2 and standard deviation 1.4 accidents. This variable is discrete. It is certainly not normaly distributed. Let x-bar be the mean number of accidents per week at the intersection during a year.(52 weeks).
Question: What is the approximate chance that there are fewer than 130 accidents at the intersection a year? (Restate in terms of x-bar)
-----------------
130 per year = 130/52 = 2.5 per week
----
(2.5-2.2)/1.4 = 1.78
So an average of 2.5 accidents perweek is 1.78 standard deviations
above the mean of 2.2.
-----
The appromate chance the average is less than 2.5 accidents per week
is about 96% or 0.9629...
====================================
Comment: Those figures are based on the distribution being normal; but
they are suitable as an approximation.
====================================
Cheers,
Stan H.